OpenStudy (anonymous):

how to find the equation of a line perpendicular to Y=3x^2-4

OpenStudy (anonymous):

It will change depending on the point to which you want it to be perpendicular.

OpenStudy (anonymous):

Suppose you want it to be perpendicular at \((x_0,y(x_0))\)

OpenStudy (anonymous):

oh yeah, at the point (2,1) sorry

OpenStudy (anonymous):

Find the instantaneous with the derivative. To make it perpendicular, you invert it and multiply it by negative one.

OpenStudy (anonymous):

So the slope is: \[ m = -\frac{1}{y'(2)} \]

OpenStudy (anonymous):

Using slope point formula: \[ (y-y_0) = m(x-x_0) \]In this case we have: \[ y-1 = \frac{1}{f'(2)}(x-2) \]

OpenStudy (anonymous):

I changed to \(f\) because in this case \(y\) is the perpendicular line, while I want \(f\) to represent our original curve.

OpenStudy (anonymous):

Perfect. Thank you!